The voltage across a 1.1-kW toaster that produces a current of 5 A is




1
Expert's answer
2021-11-04T00:41:23-0400

"P= VI\\
V= \frac{P}{I}\\
V= \frac{1.1 kW}{10A}\\
V= \frac{1100}{10}\\
V=110 v"